# Lumina[Uncaptioned image]: Detecting Hallucinations in RAG System with Context–Knowledge Signals

### v2

@@ 1 Introduction
Large language models (LLMs) are prone to hallucination, i.e., producing responses that are factually incorrect, nonsensical, or not grounded in the input or available data, while still appearing fluent and plausible (Luo et al., 2024; Huang et al., 2024; Park et al., 2025). One commonly used strategy to mitigate hallucination is providing LLMs with relevant information retrieved from external knowledge bases, so-called Retrieval-Augmented Generation (RAG) (Shuster et al., 2021; Fan et al., 2024; Gao et al., 2024). However, despite having sufficient and relevant retrieved documents, RAG systems still have a chance to hallucinate and produce statements that are either unsupported or contradict the retrieved information (Niu et al., 2024; Ridder & Schilling, 2025).
Recent work has shown that such failures often arise from conflicts between an LLM’s internal knowledge and the retrieved external context (Xu et al., 2024). In these cases, models tend to over-rely on internal knowledge regardless of correctness, undermining factual reliability (Longpre et al., 2021; Li et al., 2023; Sun et al., 2025a; Yamin et al., 2025). Inspired by this observation, recent approaches attempt to quantify hallucinations in RAG (Sun et al., 2025b; Wang, 2025; Tao et al., 2025). However, existing methods rely on mechanistic interpretability heuristics—such as selecting specific attention heads or transformer layers to achieve the optimal hallucination detection performance—which require heavy hyperparameter tuning and often fail to generalize across models and datasets.
@@ 2.1 Problem Formulation and Motivation
RAG systems aim to improve factuality by incorporating external documents into the generation process. However, hallucinations still occur when a model over-relies on its internal parametric knowledge and under-utilizes the retrieved external context. We provide a formal definition below.
@@ 2.2 Related Work
Prior works have attempted to quantify $\mathcal{E}_{p_{\theta}}(a_{t}|q,d,a_{<t})$ and $\mathcal{I}_{p_{\theta}}(a_{t}|q,d,a_{<t})$ using empirical metrics (Sun et al., 2025b; Wang, 2025). For example, Sun et al. (2025b) proposed ReDeEP, which measures external context utilization through cosine similarity between the generated token and tokens in context that have high attention weights w.r.t. certain attention heads. For internal knowledge utilization, it measures the Jensen-Shannon (JS) divergence between the hidden states before/after the FFN layer of certain transformer layers. The success of ReDeEP on some RAG hallucination detection datasets validates the idea of ?THM? LABEL:prop. Wang (2025) combine the idea of ReDeEP with semantic entropy probes (SEP) (Han et al., 2024). They quantified external context utilization by measuring the semantic correlation between the semantic entropy of the generated token and attended tokens in the context. For internal knowledge utilization, they measured the absolute difference between the semantic entropy corresponding to hidden states before and after the FFN layer.
@@ 3.1 Quantifying External Context Utilization
Definition 3.1 (Maximum Mean Discrepancy (Gretton et al., 2012)).
@@ 3.2 Quantifying Internal Knowledge Utilization
To quantify the utilization of internal knowledge, we focus on the signals in internal states of an LLM. Specifically, a transformer-based autoregressive LLM has multiple layers, through which information is gradually added into a residual stream that flows from the input layer to the output layer, shaping the output token representation and probability distribution (Geva et al., 2022). Studies have found that by projecting the hidden state of each layer to the token representation space, we can interpret what an LLM believes after the process of each layer (nostalgebraist, 2020). In addition, via logit lens (nostalgebraist, 2020), studies have identified the saturation event in an LLM, i.e., the top-$k$ prediction of the LLM remains constant in all subsequent layers after a certain layer called the $k$-th saturation layer (Geva et al., 2022; Lioubashevski et al., 2025).
@@
where $H(\cdot)$ is the entropy function, and $f$ is the logit lens  (nostalgebraist, 2020) that projects the hidden state of each layer to logits using the LayerNorm and the unembedding matrix $\bm{W}$}, i.e.,
@@ 3.3 Statistical Validation of the Measurement
In this section, we validate the soundness of our approach. Previous work such as  Sun et al. (2025b) primarily verified whether their scores have a causal relationship with hallucination but failed to show the relationship between the scores and actual external context/internal knowledge utilization. To address this, we directly assess whether our measurements capture the intended notion of utilization. Specifically, we derive verifiable implications that must hold if our proposed measurements are valid. We then use the proposed score to verify these implications with statistical hypothesis testing. If the proposed score passes all tests, the score reflects the corresponding utilization.
@@
To examine H1, we utilize data in the QA set of RAGTruth (Niu et al., 2024). We use the original data to compute $\mathcal{E}_{p_{\theta}}(a_{t}|q,d,a_{<t})$, and generate additional answers without providing retrieved documents as $a^{\prime}$ to compute $\mathcal{E}_{p_{\theta}}(a^{\prime}_{t}|q,\varnothing,a^{\prime}_{<t})$. For H2, we utilize the Summary and QA set of RAGTruth; for H4, the Summary and Data2Text set; and for H3, the entire RAGTruth dataset. We test the hypotheses with four different instruction-tuned LLMs, including Llama2-{7B, 13B} (Llama Team, 2023), Llama3-8B (Llama Team, 2024), and Mistral-7B (Jiang et al., 2023). Results in Table 1 indicate that all four implications reject their null hypothesis, validating our measurements for external context utilization and internal knowledge utilization.
@@ 4.1 Experimental Settings
We compare Lumina with baselines across 8 different hallucination detection strategies: (1) Uncertainty-based, which detects hallucination by estimating uncertainty via token-level probability or entropy. Baselines of this category include Perplexity (Ren et al., 2023), LN-Entropy (Malinin & Gales, 2021), and Focus (Zhang et al., 2023). (2) Cross-sample consistency, which detects hallucination by sampling multiple responses for a query and measuring their (logic/semantic) consistency. Approaches include SelfCKGPT (Manakul et al., 2023) and EigenScore (Chen et al., 2024). (3) Verbalization, which detects hallucinations by prompting another LLM to score the correctness of the answer. Approaches include P(True) (Kadavath et al., 2022) and RefChecker (Hu et al., 2024). (4) Utilization of external context and internal knowledge, which decouples these two signals via findings in the study of mechanistic interpretability. Baseline of this category is ReDeEP (Sun et al., 2025b). Details of each baseline are introduced in Appendix B.
@@
Experiments are conducted on two representative RAG hallucination detection benchmarks: RAGTruth (Niu et al., 2024), the first high-quality RAG hallucination detection dataset, consisting of three types of RAG tasks, including question answering, data-to-text writing, and news summarization. HalluRAG (Ridder & Schilling, 2025), a dataset of free-form question answering in an RAG setting. Details of these datasets are introduced in Appendix C.
@@
4.2 Main Results
Lumina achieves state-of-the-art performance.
Table 2 summarizes the experimental comparison across methods. The results show that Lumina has a consistently high performance across datasets and LLMs. In particular, it almost always outperforms ReDeEP, the previous attempt of measuring the utilization of external context and internal knowledge to detect hallucinations. The gap between them is particularly large on the HalluRAG dataset. Noticeably, Lumina achieves more than $0.9$ AUROC on the HalluRAG dataset across models, outperforming the baselines by a substantial margin. We further conduct an error analysis to see when and why Lumina fails. Specifically, we sample 20 false-negative and false-positive cases from the RAGTruth dataset, respectively, and qualitatively analyze the reason of errors. The result reveals that most of the errors stem from incorrect labels and low-quality retrieved documents of the dataset, suggesting a potentially higher performance in a setting with high-quality data. The details of this analysis can be found in Appendix F.
Comparison with supervised approach.
We also compare Lumina with SAPLMA (Azaria & Mitchell, 2023), a supervised approach that trained a binary classifier on the last token hidden states to detect hallucination. Since our method is unsupervised in nature and does not rely on labeled data, the supervised baseline can be viewed as a performance upper bound. Results in Appendix E.2 show that Lumina achieves a competitive performance against SAPLMA and even sometimes outperforms it, all without any training, highlighting both its supreme performance and ease of deployment.
@@ 4.3 Relaxing Assumptions
In Section 3, we implicitly make two assumptions: 1) perfect context assumption: we assume the retrieved documents $d$ are correct, sufficient, and relevant to the query. 2) same LLM assumption: we assume the LLM used to compute the external context score and internal knowledge score is the same as the LLM used to generate responses. These two assumptions are usually introduced in other hallucination detection works as well (Zhang et al., 2023; Sun et al., 2025b). Unfortunately, they are often strong and have a significant impact on the performance, limiting the usability of these methods (such as for open-sourced model-generated responses only). In this section, we investigate the performance of Lumina when relaxing these two assumptions, showing the robustness of Lumina.
@@
We relax this assumption by using different LLMs to compute the scores for a response. Specifically, we use Llama2-7B, Llama2-13B, Llama3-8B, Mistral-7B, and Qwen2.5-7B (Qwen, 2025) to detect hallucination on the RAGTruth dataset, which contains responses generated by Llama2-7B, Llama2-13B, Llama2-70B, Llama3-8B, Mistral-7B, GPT-3.5, and GPT-4. Figure 3 shows AUROC across different generator-detector LLM pairs.
The results show that the same model setting is not always necessary. Specifically, Llama2-7B achieves a comparable or higher AUROC than Llama3-8B on Llama3-8B responses. Moreover, Lumina with Llama2-7B and Llama3-8B has stable performance across different generation LLMs. Overall, Lumina demonstrates a plausible solution for generation LLM-agnostic hallucination detection, which is more practical in real-world scenarios.
@@ 4.4 Ablation Study
Impact of external context & internal knowledge.
Our final hallucination score is the combination of the external context score and internal knowledge score. To obtain more insights into how each component contributes to the final score, we ablate on the components by considering only the external context score and internal knowledge score. The right plot of Figure 4 shows that combining scores of external context and internal knowledge achieves the highest AUPRC on the RAGTruth dataset for every LLM. For example, on Llama2-13B, the combination leads to more than 10% improvement. This observation justifies the effectiveness of the hallucination score introduced in ?THM? LABEL:def:hallu_score. In addition, the left plot of Figure 4 shows that a response generated by Llama2-13B is more likely to be hallucination if it has a high internal knowledge score and a low external context score. This observation validates ?THM? LABEL:prop and suggests that Eq. (1) does not imply an objective function that forces LLM only using external context to answer questions. Instead, it suggests that the internal knowledge utilization should be grounded in an external context to achieve a reliable generation. Further experiments are shown in Appendix E.3.
@@ 5 Conclusion
This work introduce Lumina, a novel way to estimate the utilization of external context and internal knowledge when an LLM generating response with the RAG setup. Lumina significantly improves the performance of hallucination detection, which will help increase the reliability of RAG systems in real-world deployments and reduce the risk of sharing misinformation. Through a deeper analysis on Lumina in the future, researchers may better understand how LLMs utilize external context and internal knowledge to generate responses. Such findings will help the community design approaches to mitigate hallucinations and create a more reliable AI system.
@@
We provide all details of the implementation of Lumina in Appendix. D, including the approximation of MMD, the selection of kernel, and choice of random documents for measuring external context score, as well as the calibration of internal knowledge score. In Sec. 4.1, we illustrate the experimental settings, including baselines, datasets, LLMs, and evaluation metrics. The details of baselines and datasets are further provided in Appendix B and C, respectively. Furthermore, we provide the codebase of Lumina at https://anonymous.4open.science/r/LUMINA-E71B. These comprehensive reports will help future studies easily reproduce our experiments.
@@ Appendix B Details of Baselines
  • SelfCKGPT: This approach sampled multiple responses and used an NLI model to check the logistic consistency between the target generation and additional samples. In our experiment, we follow the setting of Manakul et al. (2023) to set the sample size as 20.
@@
  • RefChecker: This approach prompted an LLM to extract claims from generation, and prompted another LLM to verify the logical consistency between each claim and reference documents. In our experiment, we use dongyru/Mistral-7B-Claim-Extractor, the model finetuned by Hu et al. (2024), to extract claims.

### v3

@@ Abstract
Lumina: https://github.com/deeplearning-wisc/LUMINA
@@ 1 Introduction
Large language models (LLMs) are prone to hallucination, i.e., producing responses that are factually incorrect, nonsensical, or not grounded in the input or available data, while still appearing fluent and plausible (Luo2024HallucinationDA; 10.1145/3703155; park2025steer). One commonly used strategy to mitigate hallucination is providing LLMs with relevant information retrieved from external knowledge bases, so-called Retrieval-Augmented Generation (RAG) (shuster-etal-2021-retrieval-augmentation; 10.1145/3637528.3671470; gao2024retrievalaugmentedgenerationlargelanguage). However, despite having sufficient and relevant retrieved documents, RAG systems still have a chance to hallucinate and produce statements that are either unsupported or contradict the retrieved information (niu-etal-2024-ragtruth; ridder2025halluragdatasetdetectingcloseddomain).
Recent work has shown that such failures often arise from conflicts between an LLM’s internal knowledge and the retrieved external context (xu-etal-2024-knowledge-conflicts). In these cases, models tend to over-rely on internal knowledge regardless of correctness, undermining factual reliability (longpre-etal-2021-entity; li-etal-2023-large; sun2025seenunseendisruptiveeffect; yamin2025llmsstruggleperformcounterfactual). Inspired by this observation, recent approaches attempt to quantify hallucinations in RAG (sun2025redeep; wang2025seredeephallucinationdetectionretrievalaugmented; tao2025lostinthelaterframeworkquantifyingcontextual). However, existing methods rely on mechanistic interpretability heuristics—such as selecting specific attention heads or transformer layers to achieve the optimal hallucination detection performance—which require heavy hyperparameter tuning and often fail to generalize across models and datasets.
@@ 2.1 Problem Formulation and Motivation
RAG systems aim to improve factuality by incorporating external documents into the generation process. In cases such as news summarization, information extraction given a json file, and question answering that requires information emerging after the model’s release date, RAG is usually necessary because an LLM cannot rely solely on its internal knowledge to complete the task. However, in such cases, hallucinations still occur when a model over-relies on its internal parametric knowledge and under-utilizes the retrieved external context. We provide a formal definition below.
@@
?THM? LABEL:prop is built on a principled intuition that, if a LLM requires external knowledge to complete a task and if a retriever can provide the LLM sufficient external information, the LLM should utilize those external context and ground its reasoning ability on those context. Therefore, a response in this scenario will be considered less reliable if it disproportionally relies on the LLM’s internal knowledge without a sufficient amount of external knowledge utilization.
@@ 2.2 Related Work
Prior works have attempted to quantify $\mathcal{E}_{p_{\theta}}(a_{t}|q,d,a_{<t})$ and $\mathcal{I}_{p_{\theta}}(a_{t}|q,d,a_{<t})$ using empirical metrics (sun2025redeep; wang2025seredeephallucinationdetectionretrievalaugmented). For example, sun2025redeep proposed ReDeEP, which measures external context utilization through cosine similarity between the generated token and tokens in context that have high attention weights w.r.t. certain attention heads. For internal knowledge utilization, it measures the Jensen-Shannon (JS) divergence between the hidden states before/after the FFN layer of certain transformer layers. The success of ReDeEP on some RAG hallucination detection datasets validates the idea of ?THM? LABEL:prop. wang2025seredeephallucinationdetectionretrievalaugmented combine the idea of ReDeEP with semantic entropy probes (SEP) (han2024semantic). They quantified external context utilization by measuring the semantic correlation between the semantic entropy of the generated token and attended tokens in the context. For internal knowledge utilization, they measured the absolute difference between the semantic entropy corresponding to hidden states before and after the FFN layer.
@@ 3.1 Quantifying External Context Utilization
Definition 3.1 (Maximum Mean Discrepancy (Gretton_Borgwardt_Rasch_Schölkopf_Smola_2012)).
@@ 3.2 Quantifying Internal Knowledge Utilization
To quantify the utilization of internal knowledge, we focus on the signals in internal states of an LLM. Specifically, a transformer-based autoregressive LLM has multiple layers, through which information is gradually added into a residual stream that flows from the input layer to the output layer, shaping the output token representation and probability distribution (geva-etal-2022-transformer). Studies have found that by projecting the hidden state of each layer to the token representation space, we can interpret what an LLM believes after the process of each layer (nostalgebraist_2020). In addition, via logit lens (nostalgebraist_2020), studies have identified the saturation event in an LLM, i.e., the top-$k$ prediction of the LLM remains constant in all subsequent layers after a certain layer called the $k$-th saturation layer (geva-etal-2022-transformer; Lioubashevski_Schlank_Stanovsky_Goldstein).
@@
where $H(\cdot)$ is the entropy function, and $f$ is the logit lens  (nostalgebraist_2020) that projects the hidden state of each layer to logits using the LayerNorm and the unembedding matrix $\bm{W}$}, i.e.,
@@ 3.3 Statistical Validation of the Measurement
In this section, we validate the soundness of our approach. Previous work such as  sun2025redeep primarily verified whether their scores have a causal relationship with hallucination but failed to show the relationship between the scores and actual external context/internal knowledge utilization. To address this, we directly assess whether our measurements capture the intended notion of utilization. Specifically, we derive verifiable implications that must hold if our proposed measurements are valid. We then use the proposed score to verify these implications with statistical hypothesis testing. If the proposed score passes all tests, the score reflects the corresponding utilization.
@@
To examine H1, we utilize data in the QA set of RAGTruth (niu-etal-2024-ragtruth). We use the original data to compute $\mathcal{E}_{p_{\theta}}(a_{t}|q,d,a_{<t})$, and generate additional answers without providing retrieved documents as $a^{\prime}$ to compute $\mathcal{E}_{p_{\theta}}(a^{\prime}_{t}|q,\varnothing,a^{\prime}_{<t})$. For H2, we utilize the Summary and QA set of RAGTruth; for H4, the Summary and Data2Text set; and for H3, the entire RAGTruth dataset. We test the hypotheses with four different instruction-tuned LLMs, including Llama2-{7B, 13B} (touvron2023llama2openfoundation), Llama3-8B (grattafiori2024llama3herdmodels), and Mistral-7B (jiang2023mistral7b). Results in Table 1 indicate that all four implications reject their null hypothesis, validating our measurements for external context utilization and internal knowledge utilization.
@@ 4.1 Experimental Settings
We compare Lumina with baselines across 8 different hallucination detection strategies: (1) Uncertainty-based, which detects hallucination by estimating uncertainty via token-level probability or entropy. Baselines of this category include Perplexity (ren2023outofdistribution), LN-Entropy (malinin2021uncertainty), and Focus (zhang-etal-2023-enhancing-uncertainty). (2) Cross-sample consistency, which detects hallucination by sampling multiple responses for a query and measuring their (logic/semantic) consistency. Approaches include SelfCKGPT (manakul-etal-2023-selfcheckgpt) and EigenScore (chen2024inside). (3) Verbalization, which detects hallucinations by prompting another LLM to score the correctness of the answer. Approaches include P(True) (kadavath2022languagemodelsmostlyknow) and RefChecker (hu2024refcheckerreferencebasedfinegrainedhallucination). (4) Utilization of external context and internal knowledge, which decouples these two signals via findings in the study of mechanistic interpretability. Baseline of this category is ReDeEP (sun2025redeep). Details of each baseline are introduced in Appendix B.
@@
Experiments are conducted on two representative RAG hallucination detection benchmarks: RAGTruth (niu-etal-2024-ragtruth), the first high-quality RAG hallucination detection dataset, consisting of three types of RAG tasks, including question answering, data-to-text writing, and news summarization. HalluRAG (ridder2025halluragdatasetdetectingcloseddomain), a dataset of free-form question answering in an RAG setting. Details of these datasets are introduced in Appendix C.
@@
4.2 Main Results
Lumina achieves state-of-the-art performance.
Table 2 summarizes the experimental comparison across methods. The results show that Lumina has a consistently high performance across datasets and LLMs. In particular, it almost always outperforms ReDeEP, the previous attempt of measuring the utilization of external context and internal knowledge to detect hallucinations. The gap between them is particularly large on the HalluRAG dataset. Noticeably, Lumina achieves more than $0.9$ AUROC on the HalluRAG dataset across models, outperforming the baselines by a substantial margin. We further conduct an error analysis to see when and why Lumina fails. Specifically, we sample 20 false-negative and 20 false-positive cases from the RAGTruth dataset, respectively, and qualitatively analyze the reason of errors. The result reveals that most of the errors stem from incorrect labels and low-quality retrieved documents of the dataset, suggesting a potentially higher performance in a setting with high-quality data. The details of this analysis can be found in Appendix F.
Comparison with supervised approach.
We also compare Lumina with SAPLMA (azaria-mitchell-2023-internal), a supervised approach that trained a binary classifier on the last token hidden states to detect hallucination. Since our method is unsupervised in nature and does not rely on labeled data, the supervised baseline can be viewed as a performance upper bound. Results in Appendix E.2 show that Lumina achieves a competitive performance against SAPLMA and even sometimes outperforms it, all without any training, highlighting both its supreme performance and ease of deployment.
@@ 4.3 Relaxing Assumptions
In Section 3, we implicitly make two assumptions: 1) perfect context assumption: we assume the retrieved documents $d$ are correct, sufficient, and relevant to the query. 2) same LLM assumption: we assume the LLM used to compute the external context score and internal knowledge score is the same as the LLM used to generate responses. These two assumptions are usually introduced in other hallucination detection works as well (zhang-etal-2023-enhancing-uncertainty; sun2025redeep). Unfortunately, they are often strong and have a significant impact on the performance, limiting the usability of these methods (such as for open-sourced model-generated responses only). In this section, we investigate the performance of Lumina when relaxing these two assumptions, showing the robustness of Lumina.
@@
We relax this assumption by using different LLMs to compute the scores for a response. Specifically, we use Llama2-7B, Llama2-13B, Llama3-8B, Mistral-7B, Qwen3-4B (yang2025qwen3technicalreport), Llama4-17B (llama4) to detect hallucination on the RAGTruth dataset, which contains responses generated by Llama2-7B, Llama2-13B, Llama2-70B, Llama3-8B, Mistral-7B, GPT-3.5, and GPT-4. Figure 3 shows AUROC across different generator-detector LLM pairs.
The results show that the same model setting is not always necessary. Specifically, Llama2-7B achieves a comparable or higher AUROC than Llama3-8B on Llama3-8B responses. Moreover, Lumina with Llama2-7B and Llama3-8B has stable performance across different generation LLMs. In addition, newer models, such as Qwen3-4B and Llama4-17B, also perform well across generation LLMs. Overall, Lumina demonstrates a plausible solution for generation LLM-agnostic hallucination detection, which is more practical in real-world scenarios.
@@ 4.4 Ablation Study
Impact of external context & internal knowledge.
Our final hallucination score is the combination of the external context score and internal knowledge score. To obtain more insights into how each component contributes to the final score, we ablate on the components by considering only the external context score and internal knowledge score. The right plot of Figure 5 shows that combining scores of external context and internal knowledge achieves the highest AUPRC on the RAGTruth dataset for every LLM. For example, on Llama2-13B, the combination leads to more than 10% improvement. This observation justifies the effectiveness of the hallucination score introduced in ?THM? LABEL:def:hallu_score. In addition, the left plot of Figure 5 shows that a response generated by Llama2-13B is more likely to be hallucination if it has a high internal knowledge score and a low external context score. This observation validates ?THM? LABEL:prop and suggests that Eq. (1) does not imply an objective function that forces LLM only using external context to answer questions. Instead, it suggests that the internal knowledge utilization should be grounded in an external context to achieve a reliable generation, implying its potential for generalizing to reasoning-intensive tasks.
Additional ablations.
We also conduct other ablations, covering the selection of $\lambda$ in Eq (1), the impact of random documents $d^{\prime}$, and the contribution of two components of the information processing rate. Please see Appendix E.3 and E.4 for more details.
@@ 5 Conclusion
This work introduces Lumina, a novel way to estimate the utilization of external context and internal knowledge when an LLM generates responses with the RAG setup. Lumina significantly improves the performance of hallucination detection, which will help increase the reliability of RAG systems in real-world deployments and reduce the risk of sharing misinformation. Through a deeper analysis of Lumina in the future, researchers may better understand how LLMs utilize external context and internal knowledge to generate responses. Such findings will help the community design approaches to mitigate hallucinations and create a more reliable AI system.
@@
We provide all details of the implementation of Lumina in Appendix D, including the approximation of MMD, the selection of kernel, and the choice of random documents for measuring external context score, as well as the calibration of internal knowledge score. In Sec. 4.1, we illustrate the experimental settings, including baselines, datasets, LLMs, and evaluation metrics. The details of baselines and datasets are further provided in Appendix B and C, respectively. Furthermore, we provide the codebase of Lumina at https://github.com/deeplearning-wisc/LUMINA. These comprehensive reports will help future studies easily reproduce our experiments.
@@ Appendix B Details of Baselines
  • SelfCKGPT: This approach sampled multiple responses and used an NLI model to check the logistic consistency between the target generation and additional samples. In our experiment, we follow the setting of manakul-etal-2023-selfcheckgpt to set the sample size as 20.
@@
  • RefChecker: This approach prompted an LLM to extract claims from generation, and prompted another LLM to verify the logical consistency between each claim and reference documents. In our experiment, we use dongyru/Mistral-7B-Claim-Extractor, the model finetuned by hu2024refcheckerreferencebasedfinegrainedhallucination, to extract claims.